satellite73 (satellite73):

\[\frac{1}{2},\frac{3}{7},\frac{1}{3},\frac{5}{19},\frac{3}{14},...\] find \(a_{20}\)

OpenStudy (anonymous):

best problem posted today i miss foolformath and alchemy

OpenStudy (anonymous):

@Loser66 you are not allowed to answer

OpenStudy (loser66):

why?

OpenStudy (anonymous):

because you saw it already

OpenStudy (loser66):

I saw but I didn't understand, so???

OpenStudy (anonymous):

oh ok

OpenStudy (loser66):

hehehe... just poking you

OpenStudy (anonymous):

grrrr

OpenStudy (kinggeorge):

I don't suppose it's as simple as \(a_n=a_{n-4}a_{n-3}\)?

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

no, definitely not

OpenStudy (anonymous):

7/52 i cant read what it wants you to find

OpenStudy (anonymous):

no no 7/39

OpenStudy (nincompoop):

you've missed your shot.

OpenStudy (nincompoop):

twice

OpenStudy (nincompoop):

\[a_n = (n+1)/(n^2 + 3)\]

OpenStudy (anonymous):

not sure who you are saying but yup... that is what i did... which is how i got a6=7/39

OpenStudy (nincompoop):

a_20 = (20+1)/[(20^2)+3]

OpenStudy (nincompoop):

maybe \[a_{20} = (20+1)/[(20^2)+3]\]

OpenStudy (anonymous):

or \[\frac{21}{4003}\] if you prefer

OpenStudy (anonymous):

someone just asked this question, thought it was a good one time to bring some more math to the math portion of open study not that i mind finding the slope of a line mind you...

OpenStudy (nincompoop):

let me try a_6 a_6 = (6+1)/[(6^2)+3] a_6 = 7/39

OpenStudy (kinggeorge):

I have math questions...But people don't don't understand them :(

OpenStudy (nincompoop):

let me try if that works with 4th term a_4 = (4+1)/[(4^2)+3] a_4 = 5/19

OpenStudy (nincompoop):

@KingGeorge may I ask if there's any simpler way for vector addition

OpenStudy (anonymous):

probably myself included!

OpenStudy (anonymous):

remember alchemy? wild questions

OpenStudy (nincompoop):

I remember alchemy from 3rd world country performed by quack doctors -.-

OpenStudy (kinggeorge):

Vector addition? That's pretty darn simple in the first place. Alchemy was mostly before my time I think. Or I was just never on at the same time.

OpenStudy (anonymous):

back in the good old wild west days of openstudy, before all the rules and annoying children spamming the chats used to crash every night at around 8 or so

OpenStudy (nincompoop):

I've been here only for about a year, and I am still learning basic math